# On uniqueness in structured model learning

### v2

@@ 2.2 Neural networks
Let the parameterized classes in (7) be given by neural networks of the form in [31, Theorem 4] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (8) in Assumption 5, iv) holds true with $\beta=2q/D$ (with the networks attaining constant depth and width of order $m\log m$) and $\psi(m)=\tilde{c}m^{\frac{6q-3}{D}}$ for some constant $\tilde{c}>0$.
@@
Let the parameterized classes in (7) be given by neural networks of the form in [4, Theorem 1] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (8) in Assumption 5, iv) holds true with $\beta=q/D$ (with the networks attaining constant depth and width of order $m$) and $\psi(m)=\tilde{c}$ for some constant $\tilde{c}>0$.
Recall that a similar result as in Proposition 18 and Proposition 19 hold true for locally Lipschitz continuous activations (e.g. the ReQU activating the networks in [4, Theorem 1]) under the additional regularity Assumption 5, i) as discussed above directly after Remark 17. It remains to discuss the convergence of $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}\to\|\nabla f\|_{L^{\infty}(U)}$ as $m\to\infty$. The result in [4, Theorem 1] realizes also the simultaneous approximation of higher order derivatives at the loss of a poorer approximation rate. The work by [58] considers approximation in $W^{m,p}$-Sobolev norms for integer $m\geq 2$. Note that both results are stronger than the previously stated convergence. The works [27, 28] cover $W^{1,\infty}$-approximation by ReLU neural networks, thus, in particular inferring this type of convergence. However, a parameter estimation as stated in Assumption 5, iv) is not covered. Alternatively, e.g. for the result in [31, Theorem 4], one might eventually apply a lifting technique as outlined in Appendix A.3. This is possible in case $f$ attains higher regularity such as $W^{2,\infty}$- or $\mathcal{C}^{2}$-regularity
@@ 2.3.1 Linear case
In the first two cases welldefinedness of (15) is a consequence of [3, Theorem 6.1] (see also [3, Remark 6.2, Corollary 6.3] for the generalization to bounded Lipschitz domains). In the last case (with $\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$$\infty$ for $\hat{p}=\hat{q}$ which is important as $s_{\beta}=\infty$ is excluded in the first two cases) welldefinedness follows by $V\hookrightarrow W^{\omega,\hat{p}}(\Omega)$ and Hölder’s inequality. To achieve Assumption 4, ii) we need stricter conditions than (16) in general as outlined below. The following result, proven in Appendix B, covers Assumption 4 in the linear setup.
@@
Case 3. $|\beta|=2$: We may choose $s_{2e_{k}}=\infty$. As $\gamma>d$ there exists some constant $c_{\gamma}>0$ such that $\|\cdot\|_{C(\overline{\Omega})}\leq c_{\gamma}\|\cdot\|_{W^{1,\gamma}(\Omega)}$ yielding
@@ 3 The uniqueness problem
Before we move on to the limit problem and question of uniqueness let us justify the choice of regularization for $f_{\theta}\in W^{1,\infty}(U)^{N}$. The problem of using the $W^{1,\infty}(U)$-norm directly is that its powers are not strictly convex which is necessary for uniqueness issues later. This is overcome by the well known equivalence of the norms $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ on $W^{1,\infty}(U)$ for bounded domains $U$, which follows by [8, 6.12 A lemma of J.-L. Lions] and [8, Theorem 9.16 (Rellich–Kondrachov)]. That is, the space $W^{1,\infty}(U)$ may be strictly convexified under the equivalent norm $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ for $1<\rho<\infty$ with $|\cdot|_{W^{1,\infty}(U)}$ the seminorm in $W^{1,\infty}(U)$.
@@
Hence, we derive that $(t,\mathcal{J}_{\kappa}u^{m,l})\in U$ for $m$ sufficiently large by monotonicity of $\pi$ and $\|\hat{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(U)}\leq\|\hat{f}\|_{L^{\rho}(\mathbb{R}^{D})}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(\mathbb{R}^{D})}$. By convergence of the right hand side of (3) the terms $\|\varphi^{m}\|_{X_{\varphi}}$ and $\|u^{m}\|_{\mathcal{V}}$ are bounded due to coercivity of $\mathcal{R}_{0}$. Similarly boundedness of $\|f_{\theta^{m}}\|_{W^{1,\infty}(U)}$ follows using the norm equivalence of $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$. Boundedness of $\|u_{0}^{m}\|_{H}$ follows as $\lambda^{m}\to\infty$ as $m\to\infty$ together with boundedness of $\|u^{m}(0)\|_{H}$, which holds by boundedness of $\|u^{m}\|_{\mathcal{V}}$, and continuity of the initial condition map shown in Appendix C, II. Finally by $\lambda^{m}\to\infty$ as $m\to\infty$, coercivity of $\mathcal{D}_{\text{BC}}$, boundedness of $\gamma$ and boundedness of the $u^{m}$, also boundedness of $\|g^{m}\|_{\mathcal{B}}$ can be inferred. As a consequence of reflexivity of $X_{\varphi}^{N\times L}$, $\mathcal{V}^{N\times L},H^{N\times L},\mathcal{B}^{N\times L}$ and the fact that $W^{1,\infty}(U)^{N}$ is the dualspace of a separable space, we derive that there exist weakly convergent subsequences (w.l.o.g. the whole sequences as we will see subsequently that the limit is unique) and $\tilde{\varphi}\in X_{\varphi}^{N\times L},\tilde{u}\in\mathcal{V}^{N\times L},\tilde{u}_{0}\in H^{N\times L},\tilde{g}\in\mathcal{B}^{N\times L}$ and similarly a weak-* convergent subsequence and $\tilde{f}\in W^{1,\infty}(U)^{N}$ with $\varphi^{m}\rightharpoonup\tilde{\varphi}$, $u^{m}\rightharpoonup\tilde{u}$, $u_{0}^{m}\rightharpoonup\tilde{u}_{0}$, $g^{m}\rightharpoonup\tilde{g}$, $f_{\theta^{m}}\overset{*}{\rightharpoonup}\tilde{f}$ as $m\to\infty$ (by [8, Theorem 3.18] and Banach-Alaoglu-Bourbaki e.g. in [8, Theorem 3.16]). By weak lower semicontinuity and weak-* lower semicontinuity together with the previous considerations we derive
@@ 4 Conclusions
The results of this work provide a general framework that guarantees unique reconstructability in the limit of a practically useful all-at-once formulation in learning PDE models. This is particularly interesting because uniqueness of the quantities of interest is not given in general, but rather under certain conditions on the class of approximating functions and for certain regularization functionals. This provides an analysis-based guideline on which minimal conditions need to be ensured by practical implementations of PDE-based model learning setups in order to expect unique recovery of or regularization-minimizing solutions in the limit.
@@ A.1 Proof of Proposition 18 (Assumption 3 for neural networks)
is Lebesgue measurable for all $w\in L^{\hat{q}^{*}}(\Omega)$ which follows by standard arguments as $\mathcal{N}_{\theta}$ is continuous, $w,u(t,\cdot)$ Lebesgue measurable and measurability is preserved under integration. Employing Pettis Theorem (see [50, Theorem 1.34]) we obtain that $t\mapsto\mathcal{N}_{\theta}(t,u(t,\cdot))\in L^{\hat{q}}(\Omega)$ is Bochner measurable. Similarly as before one can show that for $u=((u_{1}^{k})_{k},\dots,(u_{N}^{k})_{k})\in(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}$ it holds for some generic $\tilde{c}>0$,
@@
again by $V_{k}\hookrightarrow L^{\hat{q}}(\Omega)$ using the isomorphism $L^{p}(0,T;L^{\hat{q}}(\Omega))^{p_{k}}\cong L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}})$ for $0\leq k\leq\kappa$. Finally, we derive by separability of $L^{\hat{q}}(\Omega)$ that $\mathcal{N}_{\theta}(u)$ is Bochner integrable (see [50, Section 1.5]) and by $p\geq q$ together with $L^{\hat{q}}(\Omega)\hookrightarrow W$ that also the Nemytskii operator $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$ is well-defined. ∎
@@ A.2 Proof of Proposition 19 (Assumption 5 for neural networks)
and the assertion of the lemma by taking the essential supremum over $z\in U$. Now for $z\in[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ an inner point of the preimage of $\left\{0\right\}$ under $\nabla\mathcal{N}_{\theta}$. it holds that $\nabla\mathcal{N}_{\theta}(z)=0$ implying (43). It remains to verify (43) for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ as the boundary $\partial[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]$ is a zeroset in $\mathbb{R}^{D}$. Following the proof of Lemma 34 we recover the estimation in (A.2). Again as $\theta^{m}\to\theta$, $\mathfrak{L}_{L-s,L-l-2}\circ\mathfrak{L}_{L-s-1}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-1$ and $\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-2$ as $m\to\infty$ and $w^{k+1}\mathfrak{L}_{1,k}(z)+b^{k+1}\neq 0$ for $1\leq k\leq L-2$ due to $\nabla\mathcal{N}_{\theta}(z)\neq 0$, for $m$ sufficiently large we end up in the smooth regime of $\sigma^{\prime}$ such that the previous arguments yield $\lim_{m\to\infty}\nabla\mathcal{N}_{\theta^{m}}(z)=\nabla\mathcal{N}_{\theta}(z)$ for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ impyling (43) and concluding the assertions of the lemma. ∎
Combining the Lemmata 33, 34 and 35 concludes the result in Proposition 19.
@@ B.1 Linear case
For the remaining part note that by [3, Theorem 6.1], [3, Corollary 6.3] and the choice of $s_{\beta}$ it holds true that the pointwise multiplication of functions is a continuous bilinear map
@@
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$, proving the weak convergence in (46). For that, let $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$. Then for a.e. $t\in(0,T)$ it holds that $D^{\beta}u_{n}(t)\in W^{\omega-|\beta|,\hat{p}}(\Omega)$ (with $W^{0,\hat{p}}(\Omega)=L^{\hat{p}}(\Omega)$) and $w^{*}(t)\in L^{\hat{q}^{*}}(\Omega)$. By [3, Theorem 6.1, Corollary 6.3] the inclusion $D^{\beta}u_{n}(t)w^{*}(t)\in L^{r_{\beta}}(\Omega)$ holds true with $\frac{\hat{p}\hat{q}}{\hat{q}-\hat{p}+\hat{p}\hat{q}}\leq r_{\beta}\leq\hat{q}^{*}$ and $r_{\beta}^{-1}\geq\frac{1}{\hat{p}}+\frac{1}{\hat{q}^{*}}-\frac{\omega-|\beta|}{d}$ (with strict inequality if $\hat{q}=1$). In particular by the requirements on $s_{\beta}$ we may choose $r_{\beta}=s_{\beta}^{*}$ (or equivalently $r_{\beta}^{*}=s_{\beta}$). Using that $D^{\beta}u_{n}\in L^{p}(0,T;W^{\omega-|\beta|,\hat{p}}(\Omega))$ and $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$ we derive
@@
for some generic constant $c>0$ for a.e. $t\in(0,T)$. Using that $\mathcal{V}\hookrightarrow C(0,T;W^{\omega,\hat{p}}(\Omega))$ by [50, Lemma 7.1] and the assumption $\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$, we derive that $u_{n}^{k}\rightharpoonup u_{n}$ in $C(0,T;W^{\omega,\hat{p}}(\Omega))$ as $k\to\infty$ which by boundedness of weakly convergent sequences in Banach spaces (see [8, Proposition 3.5 (iii)]) implies that $\|u_{n}^{k}(t)-u_{n}(t)\|_{W^{\omega,\hat{p}}(\Omega)}$ is bounded independently of $t\in(0,T)$ and $k\in\mathbb{N}$. Using uniform boundedness of $\|\varphi^{k}\|_{X_{\varphi}}$ for all $k\in\mathbb{N}$ implies that (50) and hence the integrand of (49) is majorized by the function $t\mapsto c\phi(t)\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}$ which is integrable since

### v3

@@ 2.2 Neural networks
Let the parameterized classes in (8) be given by neural networks of the form in [53, Theorem 4] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (9) in Assumption 5, iv) holds true with $\beta=2q/D$ (with the networks attaining constant depth and width of order $m\log m$) and $\psi(m)=\tilde{c}m^{\frac{6q-3}{D}}$ for some constant $\tilde{c}>0$.
@@
Let the parameterized classes in (8) be given by neural networks of the form in [11, Theorem 1] and $f\in\mathcal{C}^{q}(U)$ for some $q\geq 1$. Then (9) in Assumption 5, iv) holds true with $\beta=q/D$ (with the networks attaining constant depth and width of order $m$) and $\psi(m)=\tilde{c}$ for some constant $\tilde{c}>0$.
Recall that a similar result as in Proposition 19 and Proposition 20 holds true for locally Lipschitz continuous activations (e.g. the ReQU activating the networks in [11, Theorem 1]) under the additional regularity Assumption 5, i) as discussed above directly after Remark 18. It remains to discuss the convergence of $\|\nabla f_{\theta^{m}}\|_{L^{\infty}(U)}\to\|\nabla f\|_{L^{\infty}(U)}$ as $m\to\infty$. The result in [11, Theorem 1] realizes also the simultaneous approximation of higher order derivatives at the loss of a poorer approximation rate. The work by [112] considers approximation in $W^{m,p}$-Sobolev norms for integer $m\geq 2$. Note that both results are stronger than the previously stated convergence. The works [48, 49] cover $W^{1,\infty}$-approximation by ReLU neural networks, thus, in particular inferring this type of convergence. However, a parameter estimation as stated in Assumption 5, iv) is not covered. Alternatively, e.g. for the result in [53, Theorem 4], one might eventually apply a lifting technique as outlined in Appendix A.3. This is possible in case $f$ attains higher regularity such as $W^{2,\infty}$- or $\mathcal{C}^{2}$-regularity
@@ 2.3.1 Linear case
In the first two cases welldefinedness of (16) is a consequence of [9, Theorem 6.1] (see also [9, Remark 6.2, Corollary 6.3] for the generalization to bounded Lipschitz domains). In the last case (with $\frac{\hat{p}\hat{q}}{\hat{p}-\hat{q}}$$\infty$ for $\hat{p}=\hat{q}$ which is important as $s_{\beta}=\infty$ is excluded in the first two cases) welldefinedness follows by $V\hookrightarrow W^{\omega,\hat{p}}(\Omega)$ and Hölder’s inequality. To achieve Assumption 4, ii) we need stricter conditions than (17) in general as outlined below. The following result, proven in Appendix B, covers Assumption 4 in the linear setup.
@@
Case 3. $|\beta|=2$: We may choose $s_{2e_{k}}=\infty$. As $\gamma>d$ there exists some constant $c_{\gamma}>0$ such that $\|\cdot\|_{\mathcal{C}(\overline{\Omega})}\leq c_{\gamma}\|\cdot\|_{W^{1,\gamma}(\Omega)}$ yielding
@@ 3 The uniqueness problem
Before we move on to the limit problem and question of uniqueness let us justify the choice of regularization for $f_{\theta}\in W^{1,\infty}(U)^{N}$. The problem of using the $W^{1,\infty}(U)$-norm directly is that its powers are not strictly convex which is necessary for uniqueness issues later. This is overcome by the well known equivalence of the norms $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ on $W^{1,\infty}(U)$ for bounded domains $U$, which follows by [15, 6.12 A lemma of J.-L. Lions] and [15, Theorem 9.16 (Rellich–Kondrachov)]. That is, the space $W^{1,\infty}(U)$ may be strictly convexified under the equivalent norm $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$ for $1<\rho<\infty$ with $|\cdot|_{W^{1,\infty}(U)}$ the seminorm in $W^{1,\infty}(U)$.
@@
Hence, we derive that $(t,\mathcal{J}_{\kappa}u^{m,l})\in U$ for $m$ sufficiently large by monotonicity of $\pi$ and $\|\hat{f}\|_{L^{\rho}(U)}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(U)}\leq\|\hat{f}\|_{L^{\rho}(\mathbb{R}^{D})}^{\rho}+\|\nabla\hat{f}\|_{L^{\infty}(\mathbb{R}^{D})}$. By convergence of the right hand side of (3) the terms $\|\varphi^{m}\|_{X_{\varphi}}$ and $\|u^{m}\|_{\mathcal{V}}$ are bounded due to coercivity of $\mathcal{R}_{0}$. Similarly boundedness of $\|f_{\theta^{m}}\|_{W^{1,\infty}(U)}$ follows using the norm equivalence of $\|\cdot\|_{W^{1,\infty}(U)}$ and $\|\cdot\|_{L^{\rho}(U)}+|\cdot|_{W^{1,\infty}(U)}$. Boundedness of $\|u_{0}^{m}\|_{H}$ follows as $\lambda^{m}\to\infty$ as $m\to\infty$ together with boundedness of $\|u^{m}(0)\|_{H}$, which holds by boundedness of $\|u^{m}\|_{\mathcal{V}}$, and continuity of the initial condition map shown in Appendix C, II. Finally by $\lambda^{m}\to\infty$ as $m\to\infty$, coercivity of $\mathcal{D}_{\text{BC}}$, boundedness of $\gamma$ and boundedness of the $u^{m}$, also boundedness of $\|g^{m}\|_{\mathcal{B}}$ can be inferred. As a consequence of reflexivity of $X_{\varphi}^{N\times L}$, $\mathcal{V}^{N\times L},H^{N\times L},\mathcal{B}^{N\times L}$ and the fact that $W^{1,\infty}(U)^{N}$ is the dualspace of a separable space, we derive that there exist weakly convergent subsequences (w.l.o.g. the whole sequences as we will see subsequently that the limit is unique) and $\tilde{\varphi}\in X_{\varphi}^{N\times L},\tilde{u}\in\mathcal{V}^{N\times L},\tilde{u}_{0}\in H^{N\times L},\tilde{g}\in\mathcal{B}^{N\times L}$ and similarly a weak-* convergent subsequence and $\tilde{f}\in W^{1,\infty}(U)^{N}$ with $\varphi^{m}\rightharpoonup\tilde{\varphi}$, $u^{m}\rightharpoonup\tilde{u}$, $u_{0}^{m}\rightharpoonup\tilde{u}_{0}$, $g^{m}\rightharpoonup\tilde{g}$, $f_{\theta^{m}}\overset{*}{\rightharpoonup}\tilde{f}$ as $m\to\infty$ (by [15, Theorem 3.18] and Banach-Alaoglu-Bourbaki e.g. in [15, Theorem 3.16]). By weak lower semicontinuity and weak-* lower semicontinuity together with the previous considerations we derive
@@ 4 Conclusions
The results of this work provide a general framework that guarantees unique reconstructability in the limit of a practically useful all-at-once formulation in learning PDE models. This is particularly interesting because uniqueness of the quantities of interest is not given in general, but rather under certain conditions on the class of approximating functions and for certain regularization functionals. This provides an analysis-based guideline on which minimal conditions need to be ensured by practical implementations of PDE-based model learning setups in order to expect unique recovery of regularization-minimizing solutions in the limit.
@@ A.1 Proof of Proposition 19 (Assumption 3 for neural networks)
is Lebesgue measurable for all $w\in L^{\hat{q}^{*}}(\Omega)$ which follows by standard arguments as $\mathcal{N}_{\theta}$ is continuous, $w,u(t,\cdot)$ Lebesgue measurable and measurability is preserved under integration. Employing Pettis Theorem (see [98, Theorem 1.34]) we obtain that $t\mapsto\mathcal{N}_{\theta}(t,u(t,\cdot))\in L^{\hat{q}}(\Omega)$ is Bochner measurable. Similarly as before one can show that for $u=((u_{1}^{k})_{k},\dots,(u_{N}^{k})_{k})\in(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}$ it holds for some generic $\tilde{c}>0$,
@@
again by $V_{k}\hookrightarrow L^{\hat{q}}(\Omega)$ using the isomorphism $L^{p}(0,T;L^{\hat{q}}(\Omega))^{p_{k}}\cong L^{p}(0,T;L^{\hat{q}}(\Omega)^{p_{k}})$ for $0\leq k\leq\kappa$. Finally, we derive by separability of $L^{\hat{q}}(\Omega)$ that $\mathcal{N}_{\theta}(u)$ is Bochner integrable (see [98, Section 1.5]) and by $p\geq q$ together with $L^{\hat{q}}(\Omega)\hookrightarrow W$ that also the Nemytskii operator $\mathcal{N}_{\theta}:(\otimes_{k=0}^{\kappa}\mathcal{V}_{k}^{\times})^{N}\to\mathcal{W}$ is well-defined. ∎
@@ A.2 Proof of Proposition 20 (Assumption 5 for neural networks)
and the assertion of the lemma by taking the essential supremum over $z\in U$. Now for $z\in[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ an inner point of the preimage of $\left\{0\right\}$ under $\nabla\mathcal{N}_{\theta}$. it holds that $\nabla\mathcal{N}_{\theta}(z)=0$ implying (44). It remains to verify (44) for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ as the boundary $\partial[(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]$ is a zeroset in $\mathbb{R}^{D}$. Following the proof of Lemma 35 we recover the estimation in (A.2). Again as $\theta^{m}\to\theta$, $\mathfrak{L}_{L-s,L-l-2}\circ\mathfrak{L}_{L-s-1}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-1$ and $\mathfrak{L}_{L-s+1,L-l-2}\circ\mathfrak{L}_{L-s}^{m}\to\mathfrak{L}_{1,L-l-2}$ in $L^{\infty}(U)$ for $0\leq l\leq s-2$ as $m\to\infty$ and $w^{k+1}\mathfrak{L}_{1,k}(z)+b^{k+1}\neq 0$ for $1\leq k\leq L-2$ due to $\nabla\mathcal{N}_{\theta}(z)\neq 0$, for $m$ sufficiently large we end up in the smooth regime of $\sigma^{\prime}$ such that the previous arguments yield $\lim_{m\to\infty}\nabla\mathcal{N}_{\theta^{m}}(z)=\nabla\mathcal{N}_{\theta}(z)$ for $z\in[U\backslash(\nabla\mathcal{N}_{\theta})^{-1}(\left\{0\right\})]^{\circ}$ implying (44) and concluding the assertions of the lemma. ∎
Combining the Lemmata 34, 35 and 36 concludes the result in Proposition 20.
@@ B.1 Linear case
For the remaining part note that by [9, Theorem 6.1], [9, Corollary 6.3] and the choice of $s_{\beta}$ it holds true that the pointwise multiplication of functions is a continuous bilinear map
@@
in $L^{q}(0,T;L^{\hat{q}}(\Omega))$ as $k\to\infty$, proving the weak convergence in (47). For that, let $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$. Then for a.e. $t\in(0,T)$ it holds that $D^{\beta}u_{n}(t)\in W^{\omega-|\beta|,\hat{p}}(\Omega)$ (with $W^{0,\hat{p}}(\Omega)=L^{\hat{p}}(\Omega)$) and $w^{*}(t)\in L^{\hat{q}^{*}}(\Omega)$. By [9, Theorem 6.1, Corollary 6.3] the inclusion $D^{\beta}u_{n}(t)w^{*}(t)\in L^{r_{\beta}}(\Omega)$ holds true with $\frac{\hat{p}\hat{q}}{\hat{q}-\hat{p}+\hat{p}\hat{q}}\leq r_{\beta}\leq\hat{q}^{*}$ and $r_{\beta}^{-1}\geq\frac{1}{\hat{p}}+\frac{1}{\hat{q}^{*}}-\frac{\omega-|\beta|}{d}$ (with strict inequality if $\hat{q}=1$). In particular by the requirements on $s_{\beta}$ we may choose $r_{\beta}=s_{\beta}^{*}$ (or equivalently $r_{\beta}^{*}=s_{\beta}$). Using that $D^{\beta}u_{n}\in L^{p}(0,T;W^{\omega-|\beta|,\hat{p}}(\Omega))$ and $w^{*}\in L^{q^{*}}(0,T;L^{\hat{q}^{*}}(\Omega))$ we derive
@@
for some generic constant $c>0$ for a.e. $t\in(0,T)$. Using that $\mathcal{V}\hookrightarrow\mathcal{C}(0,T;W^{\omega,\hat{p}}(\Omega))$ by [98, Lemma 7.1] and the assumption $\tilde{V}\hookrightarrow W^{\omega,\hat{p}}(\Omega)$, we derive that $u_{n}^{k}\rightharpoonup u_{n}$ in $\mathcal{C}(0,T;W^{\omega,\hat{p}}(\Omega))$ as $k\to\infty$ which by boundedness of weakly convergent sequences in Banach spaces (see [15, Proposition 3.5 (iii)]) implies that $\|u_{n}^{k}(t)-u_{n}(t)\|_{W^{\omega,\hat{p}}(\Omega)}$ is bounded independently of $t\in(0,T)$ and $k\in\mathbb{N}$. Using uniform boundedness of $\|\varphi^{k}\|_{X_{\varphi}}$ for all $k\in\mathbb{N}$ implies that (51) and hence the integrand of (50) is majorized by the function $t\mapsto c\phi(t)\|w^{*}(t)\|_{L^{\hat{q}^{*}}(\Omega)}$ which is integrable since